# Aligning Attention Distribution to Information Flow for Hallucination Mitigation in Large Vision-Language Models

### v1

@@ Abstract
Due to the unidirectional masking mechanism, Decoder-Only models propagate information from left to right. LVLMs (Large Vision-Language Models) follow the same architecture, with visual information gradually integrated into semantic representations during forward propagation. Through systematic analysis, we observe that over 80% of the visual information is absorbed into the semantic representations. However, the model’s attention still predominantly focuses on the visual representations. This misalignment between the attention distribution and the actual information flow undermines the model’s visual understanding ability and contributes to hallucinations. To address this issue, we enhance the model’s visual understanding by leveraging the core information embedded in semantic representations. Specifically, we identify attention heads that focus on core semantic representations based on their attention distributions. Then, through a two-stage optimization paradigm, we propagate the advantages of these attention heads across the entire model, aligning the attention distribution with the actual information flow. We evaluate our method on three image captioning benchmarks using five different LVLMs, demonstrating its effectiveness in significantly reducing hallucinations. Further experiments reveal a trade-off between reduced hallucinations and richer details. Notably, our method allows for manual adjustment of the model’s conservativeness, enabling flexible control to meet diverse real-world requirements. Code will be released once accepted.
Aligning Attention Distribution to Information Flow for Hallucination Mitigation in Large Vision-Language Models
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: fengchong@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
@@ 1 Introduction
Large Vision-Language Models (LVLMs) Liu et al. (2024c); Dai et al. (2023); Bai et al. (2023) integrate Large Language Models (LLMs) with visual encoders, aligning the extracted visual features with the semantic space of the LLMs to enable comprehensive understanding of visual content. However, they often suffer from the problem of hallucination Ji et al. (2023); Liu et al. (2024b), where the models may generate content that is inconsistent with the visual evidence, thereby severely undermining their reliability in real-world scenarios.
@@
Recent studies have explored methods to mitigate hallucinations in LVLMs. Some works attribute hallucinations to the influence of language priors Leng et al. (2024); Zhang et al. (2025), proposing contrastive decoding-based methods to suppress language priors. However, they do not explicitly enhance the model’s visual understanding capabilities. Other studies argue that LVLMs exhibit biased mechanisms in visual attention distribution Yin et al. (2025); Kang et al. (2025); He et al. (2024), and propose optimizations such as increasing the relative weight of key visual tokens. In this work, however, we point out that these approaches make suboptimal adjustments to the information flow, as they overlook the influence of information aggregation.
@@
In fact, the unidirectional masked generation process in Transformer-based models can be regarded as a form of information flow Vaswani et al. (2017), where information from earlier tokens in the input or generated sequence flows toward later tokens. Since LVLMs adopt this architecture, their understanding of visual content can be modeled as a process in which visual information flows into the semantic representations (since visual tokens are positioned before the textual tokens). Based on this perspective, we follow the Attention Rollout Abnar and Zuidema (2020) to conduct an in-depth analysis of information flow within LVLMs. The visualization results are shown in Fig. 1.
[…]
@@
As shown in the visualization, visual information progressively flows into semantic representations across layers. In the final layers, more than 80% of the visual information is integrated into the semantic representations. This indicates that, during the forward pass of the LVLMs, visual information is gradually encoded into the semantic representations. While prior works largely concentrate on optimizing attention allocation to visual features, they neglect the critical insight that essential information has already been integrated into semantic representations, where attention refinement may be more impactful.
[…]
@@
Building upon the above findings, as the number of layers increases, textual representations progressively integrate more visual information and thus play an increasingly important role. Consequently, LVLMs should place greater attention on these integrated representations. To investigate this, we statistically analyze the attention allocation between semantic representations and others at each layer, as illustrated in Fig. 3. The model’s attention to semantic representations is significantly lower than expected, accounting for only about one-fifth of the total attention. We attribute this phenomenon to the nature of supervision in LVLM training, which is primarily driven by soft supervision from end-to-end data and lacks explicit guidance on attention allocation. Therefore, we argue that the misalignment between the attention distribution and the visual information flow is a major contributing factor to hallucinations, as the model does not sufficiently attend to regions enriched with visual information.
@@
To address the aforementioned challenge, we propose SEVI (Semantic-Enhanced Visual Interpretation), a novel training-free approach that augments the model’s attention to those semantic representations that have absorbed visual information. Furtherly, we encourage the model to focus on the most meaningful representations in order to reduce the occurrence of hallucinations. Several studies Wang et al. (2023); Xiao et al. (2024) have revealed that contextual information tends to spontaneously concentrate in some tokens, forming anchor tokens that encapsulate key information. These tokens are typically characterized by receiving disproportionately high attention, as shown in Fig. 4. Therefore, our method focuses the model’s attention on these core semantic representations at higher layers where information integration and aggregation occur. The key visual information contained within these representations provides more effective guidance for the model’s understanding of visual inputs, thereby mitigating hallucination.
@@
Figure 3: Attention distribution between textual tokens and all other tokens, obtained by averaging across all attention heads.
@@
Specifically, we identify superior attention heads within the multi-head attention mechanism—those that attend to core semantic representations—and use their attention distribution as the target distribution to guide the optimization of the overall model attention. To achieve this, we design a two-stage attention refinement paradigm, incorporating a smoothing mechanism to ensure stable performance improvements.
@@
We evaluate the effectiveness of our method in mitigating hallucinations by testing it with five LVLMs: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen2-VL-Instruct Wang et al. (2024c), and Qwen2.5-VL-Instruct Bai et al. (2025), on three image captioning benchmarks: CHAIR Rohrbach et al. (2018), AMBER Wang et al. (2024b), and DetailCaps Ye et al. (2025). Experimental results show that our approach significantly reduces hallucinations while supporting manual adjustment of the model’s conservativeness, enabling a flexible trade-off between cautious outputs (fewer hallucinations) and comprehensive descriptions (richer details).
Figure 4: Semantic attention distribution of LLaVA-1.5 during image description. We compute the proportion of peak attention within the semantic representations relative to the entire set of textual tokens.
@@
  • We reveal that visual information is indeed integrated into semantic representations. However, the model’s attention allocation does not align with this flow pattern, as it continues to focus more on visual representations.
@@ 2 Related Work
Various approaches have been proposed to mitigate hallucinations in LVLMs, such as improving training data quality Liu et al. (2024a); Sun et al. (2024); Gunjal et al. (2024) or designing specific data formats Chen et al. (2025); Hu et al. (2023); Zhai et al. (2024) to enhance model reliability. However, these training-based methods often suffer from limited scalability.
@@
In recent years, the rapid evolution of LVLM backbones has brought increasing attention to training-free methods. Early approaches Yin et al. (2024); Wang et al. (2024a); Zhou et al. (2024) primarily relied on post-processing techniques to correct hallucinated content after generation. However, the complexity of such pipelines and their dependence on external modules limited their practical usability. Researchers have since shifted focus toward investigating the root causes of hallucinations, exploring targeted optimizations in both the decoding strategies and the decoding processes. Some methods Leng et al. (2024); Wang et al. (2024d); Zhang et al. (2025); Zhao et al. (2025) apply contrastive decoding algorithms to refine the logits, thereby reducing the likelihood of hallucination-related tokens. While these approaches improve the expressiveness of the model and enhance the transfer of visual information into textual form, they do not strengthen the model’s intrinsic cross-modal understanding. Other works have proposed enhancements to the attention mechanism, such as correcting visual positional bias Li et al. (2025); Zhu et al. (2025), directly boosting visual attention Zhu et al. (2024); He et al. (2024), or redistributing attention weights Kang et al. (2025); Yin et al. (2025). However, these methods primarily concentrate on visual attention and overlook the vision-integrated semantic representation.
@@
In contrast to these approaches, we analyze the information flow between the visual and textual modalities in LVLMs and focus the model’s attention on key information within semantic representations, thereby reducing hallucinations more effectively.
3 Preliminary
@@
LVLMs consist of a visual encoder and an LLM backbone. High-dimensional features extracted by the visual encoder are concatenated with textual embeddings and jointly fed into the LLM. Subsequently, the visual and semantic features undergo cross-modal interaction through the Self-Attention layers within the LLM. The cross-modal interaction in layer $l$ simplified as $\Delta X^{l}=\text{softmax}(W)X^{l}$, where $X^{l}=[X^{l}_{V}:X^{l}_{T}]$ denotes the hidden states of visual features $X^{l}_{V}$ and semantic features $X^{l}_{T}$. $W\in\mathbb{R}^{H\times L\times L}$ is the attention weights, $H$ is the number of attention head and $L$ is the model’s context length.
@@
When generating the ($j$+1)-th token, the LVLM’s attention to feature $X^{l}_{i}$ at layer $l$ is quantified by the weight value $W_{i,j}$. A larger $W_{i,j}$ indicates that the model’s current layer relies more heavily on the information from $X^{l}_{i}$ when determining the generated content. Therefore, we can intuitively infer that more important features $X^{l}_{i}$ should generally correspond to larger attention weights $W_{i,j}$.
@@ 4 Method
Figure 5: The diagram of our method. We first categorize attention heads based on their focus on semantic representations into semantic heads and other heads. Semantic heads are further divided into core semantic heads and global semantic heads, depending on whether they attend to core semantic representations. We then align the model’s attention distribution with the core semantic heads through a two-stage optimization process.
@@
We attribute the occurrence of hallucinations to a misalignment between the attention allocation of LVLMs and the underlying information flow. To address this, we enhance the visual interpretation capabilities of LVLMs by leveraging semantic representations that encapsulate visual information. We first identify superior attention heads that focus on core semantic representations, and then propagate their strengths to other heads through a two-stage optimization paradigm, as illustrated in Fig. 5.
4.1 Attention allocation Alignment
@@
Considering that visual information is progressively integrated into semantic representations through forward propagation—and that only a few of these representations aggregate key information—an intuitive approach is to increase the model’s attention to these core semantic representations in the later layers. However, rigidly modifying the model’s attention distribution may disrupt its inherent latent features, potentially resulting in suboptimal performance.
@@
In multi-head attention mechanisms, attention heads naturally diversify after training, forming a heterogeneous and specialized structure Vaswani et al. (2017). Accordingly, we identify superior attention heads that attend to core semantic representations and transfer their attention patterns to other heads, leading to a global refinement of the model’s attention distribution. Specifically, at each layer, we calculate the attention weight of each head with respect to the semantic representations. We then identify those heads that allocate more than 50% of their attention to semantic representations:
$h\in\left\{\begin{aligned} H_{S}\quad&\text{ if }\sum W^{h}_{S}>\sum W^{h}_{O}\\ H_{O}\quad&\text{ if }\sum W^{h}_{S}\leq\sum W^{h}_{O}\end{aligned}\right.$ (1)
@@
$H_{S}$ denotes the set of heads that focus more on semantic representations, whereas $H_{O}$ denotes other heads. $H_{S}$ are further categorized based on their attention distributions into Core Semantic Heads ($H_{S_{c}}$), which focus on core semantic representations, and Global Semantic Heads ($H_{S_{g}}$), which fail to effectively attend to them:
@@
$h\in\left\{\begin{aligned} H_{S_{c}}\quad&\text{if}\quad\text{max}\overline{W}^{h}_{S}>\kappa\\ H_{S_{g}}\quad&\text{if}\quad\text{max}\overline{W}^{h}_{S}\leq\kappa\end{aligned}\right.$ (2)
@@
where $\kappa$ is a hyperparameter. We then take core semantic heads to construct a target distribution to optimize the model’s attention allocation.
@@
The model’s internal attention is then optimized through a two-stage refinement process. During this process, we construct the guidance across attention heads by performing a weighted summation of their attention weights, and incorporate smoothing to enhance the stability of feature representations.
$\hat{W}^{h}=\frac{W^{h}+\omega W^{\prime}}{1+\omega}$ (3)
where $\hat{W}^{h}$ denotes the optimized attention weights of head $h$, $W^{\prime}$ represents the target weight distribution derived from superior heads, and $\omega$ is a hyperparameter that controls the extent of the modification.
We first use global semantic heads to guide other heads, aiming to align the model’s attention distribution with the flow of visual information. To this end, we apply average pooling across global semantic heads to construct the target attention distribution. Subsequently, we use core semantic heads to guide the global semantic heads, with the goal of enhancing the model’s focus on critical information. Here, we employ max pooling across core semantic heads to form the corresponding target distribution.
$\hat{W}^{h}_{O}=\frac{W^{h}_{O}+\omega\cdot\text{avg}W^{h}_{S_{g}}}{1+\omega},$ (4)
$\hat{W}^{h}_{S_{g}}=\frac{W^{h}_{S_{g}}+\omega\cdot\text{max}W^{h}_{S_{c}}}{1+\omega}$
MaxLen Method LLaVA-1.5 InstructBLIP LLaVA-1.5-13B
@@
64 Regular 24.4 8.9 56.6 35.6 13.2 56.4 24.0 7.8 56.5
VCD 25.0 8.3 59.0 32.2 10.3 60.6 23.4 7.5 59.4
ICD 23.2 8.1 58.4 29.8 9.8 60.6 18.8 6.6 58.4
IBD 21.2 6.9 58.8 27.8 9.2 60.3 22.4 7.1 59.7
VAR 25.2 8.6 55.3 - - - 25.8 8.5 56.3
VAF 26.2 9.3 56.6 32.0 12.0 55.3 23.2 8.0 57.6
DeFG 22.4 7.2 58.2 32.4 11.0 59.4 22.4 6.9 59.6
CICD 18.0 6.1 59.6 23.8 7.7 62.2 20.4 6.6 59.4
@@
512 Regular 54.6 16.4 72.6 62.6 19.5 66.9 58.8 17.0 73.4
VCD 59.8 17.8 75.6 64.8 18.8 71.9 60.2 16.4 76.9
ICD 57.0 15.0 74.6 59.0 17.1 69.2 55.0 14.5 76.4
IBD 57.6 16.5 74.2 57.6 15.7 70.8 50.6 14.3 76.1
VAR 60.0 18.3 72.6 - - - 54.8 15.3 73.8
VAF 58.8 19.0 71.3 58.6 17.8 66.9 57.2 16.2 73.7
DeFG 57.4 16.3 75.5 59.0 17.7 71.5 51.8 14.2 75.5
CICD 43.8 11.7 75.0 49.8 13.7 70.3 44.4 12.9 76.0
@@
Table 1: Results on CHAIR. SEVIBalanced and SEVIFocused represent our method operating in the Balanced and Focused mode, respectively. MaxLen denotes the maximum generation length.

### v3

# Mitigating Hallucination in Large Vision-Language Models through Aligning Attention Distribution to Information Flow

@@ Abstract
Due to the unidirectional masking mechanism, Decoder-Only models propagate information from left to right. LVLMs (Large Vision-Language Models) follow the same architecture, with visual information gradually integrated into semantic representations during forward propagation. Through systematic analysis, we observe that the majority of the visual information is absorbed into the semantic representations. However, the model’s attention distribution does not exhibit sufficient emphasis on semantic representations. This misalignment between the attention distribution and the actual information flow undermines the model’s visual understanding ability and contributes to hallucinations. To address this issue, we enhance the model’s visual understanding by leveraging the core information embedded in semantic representations. Specifically, we identify attention heads that focus on core semantic representations based on their attention distributions. Then, through a two-stage optimization paradigm, we propagate the advantages of these attention heads across the entire model, aligning the attention distribution with the actual information flow. We evaluate our method on three image captioning benchmarks using five different LVLMs, demonstrating its effectiveness in significantly reducing hallucinations. Further experiments reveal a trade-off between reduced hallucinations and richer details. Notably, our method allows for manual adjustment of the model’s conservativeness, enabling flexible control to meet diverse real-world requirements. [^1] Code is available at https://github.com/beta-nlp/SEVI.
Mitigating Hallucination in Large Vision-Language Models through Aligning Attention Distribution to Information Flow
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1,11footnotemark: 1, Chong Feng1,3,thanks: Corresponding Authors. 1School of Computer Science and Technology, Beijing Institute of Technology 2Zhongguancun Academy 3Southeast Academy of Information Technology, Beijing Institute of Technology {zhqingan, bit_zhangfeng, sunxin, fengchong}@bit.edu.cn
@@ 1 Introduction
Large Vision-Language Models (LVLMs) Dai et al. (2023); Bai et al. (2023); Liu et al. (2024c) integrate Large Language Models (LLMs) with visual encoders, aligning the extracted visual features with the semantic space of the LLMs to enable comprehensive understanding of visual content. However, they often suffer from hallucination Wang et al. (2023b); Li et al. (2023); Liu et al. (2024b), where the models may generate content that is inconsistent with the visual evidence, thereby severely limiting their reliability in real-world scenarios.
@@
Recent studies have explored methods to mitigate hallucinations in LVLMs. Leng et al. (2024) and Zhang et al. (2025) attribute hallucinations to the influence of language priors, proposing contrastive decoding-based methods to suppress language priors. However, they do not explicitly enhance the model’s visual understanding capabilities. Other studies Yin et al. (2025); Kang et al. (2025); He et al. (2025) argue that LVLMs exhibit biased mechanisms in visual attention distribution, and propose optimizations such as increasing the relative weight of key visual tokens. In this work, however, we point out that these approaches make suboptimal adjustments to the information flow, as they overlook the influence of information aggregation.
@@
In fact, the unidirectional masked generation process in Transformer-based models can be regarded as a form of information flow, where information from earlier tokens in the input or generated sequence flows toward later tokens. Since LVLMs adopt this architecture, their understanding of visual content can be modeled as a process in which visual information flows into the semantic representations (since visual tokens are positioned before the textual tokens). Based on this perspective, we follow the Attention Rollout Abnar and Zuidema (2020) to conduct an in-depth analysis of information flow within LVLMs. The visualization results are shown in Fig. 1.
Specifically, we use $\mathbf{F}_{i,j}^{l}$ to represent the contributions of the $i\text{-th}$ input embedding to the $j\text{-th}$ representation in layer $l$. Considering the residual connection in the forward propagation, we formulate the recursive relation as $\mathbf{F}^{l}=(\mathbf{I}+\mathbf{W}_{\mathrm{attn}})/2\cdot\mathbf{F}^{l-1}$, where $\mathbf{W}_{\mathrm{attn}}$ is the attention weight after average pooling across multi-heads. The contributions of each input embedding were normalized to 1 across all representations. We perform layer-wise quantification of visual embeddings’ contributions to the visual and semantic representations, respectively.
@@
As shown in the visualization, visual information progressively flows into semantic representations across layers. In the final layers, the majority of the visual information is integrated into the semantic representations. This indicates that during the forward pass of the LVLMs, visual information is gradually encoded into the semantic representations. While prior works largely concentrate on optimizing the attention distribution to visual features, they neglect the critical insight that essential information has already been integrated into semantic representations, where attention refinement may be more impactful.
Figure 2: Jensen-Shannon divergence (JSD) between the regular and vision-masked logits distribution. layer=l represents the application of masking vision features starting from the $l\text{-th}$ layer. Lower $\log(\text{JSD})$ denotes higher consistency between two distributions. The generation context is The image depicts a _.
@@
Building upon the above findings, as the number of layers increases, textual representations progressively integrate more visual information and thus play an increasingly important role. Consequently, LVLMs should place greater attention on these integrated representations. To investigate this, we statistically analyze the attention allocation between semantic representations and others at each layer, as illustrated in Fig. 3. The model’s attention to semantic representations is significantly lower than expected, accounting for only about one-fifth of the total attention. This phenomenon stems from the nature of supervision in LVLM training, which is primarily driven by soft supervision from end-to-end data and lacks explicit guidance on the attention distribution. We argue that the model’s attention distribution should align with the actual flow pattern of visual information. Otherwise, hallucinations may arise, as the model does not sufficiently attend to regions enriched with visual information.
@@
To address the aforementioned challenge, we propose SEVI (Semantic-Enhanced Visual Interpretation), a novel training-free approach that augments the model’s attention to those semantic representations that have absorbed visual information. Furtherly, we encourage the model to focus on the most meaningful representations in order to reduce the occurrence of hallucinations. Several studies Wang et al. (2023c); Xiao et al. (2024) have revealed that contextual information tends to spontaneously concentrate in some tokens, forming anchor tokens that encapsulate core information. These tokens are typically characterized by receiving disproportionately high attention Wang et al. (2023c). Therefore, our method focuses the model’s attention on these core semantic representations at higher layers where information integration and aggregation occur. The core visual information contained within these representations provides more effective guidance for the model’s understanding of visual inputs, thereby mitigating hallucination.
@@
Figure 3: Attention distribution between semantic representations and all others, obtained by averaging across all attention heads.
@@
Specifically, we identify superior attention heads within the multi-head attention mechanism—those that attend to core semantic representations—and use their attention distribution as the target distribution to guide the optimization of the model’s overall attention. To achieve this, we design a two-stage attention optimization paradigm, incorporating a smoothing mechanism to ensure stable performance improvements.
@@
We evaluate the effectiveness of our method in mitigating hallucinations by testing it with five LVLMs: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025), on three image captioning benchmarks: CHAIR Rohrbach et al. (2018), AMBER Wang et al. (2023a), and DetailCaps Ye et al. (2025). Experimental results show that our approach significantly reduces hallucinations while supporting manual adjustment of the model’s conservativeness, enabling a flexible trade-off between cautious outputs (fewer hallucinations) and comprehensive descriptions (richer details).
@@
  • We reveal that visual information is indeed integrated into semantic representations. However, the model’s attention distribution does not align with this flow pattern, leading to hallucinations.
@@ 2 Related Work
Various approaches have been proposed to mitigate hallucinations in LVLMs, such as improving training data quality Liu et al. (2024a); Sun et al. (2024); Gunjal et al. (2024) or designing specific data formats Hu et al. (2023); Zhai et al. (2024); Chen et al. (2025) to enhance model reliability. However, these training-based methods often suffer from limited scalability.
@@
In recent years, the rapid evolution of LVLM backbones has brought increasing attention to training-free methods. Early approaches Yin et al. (2024); Wang et al. (2024a); Zhou et al. (2024) primarily relied on post-processing techniques to correct hallucinated content after generation. However, the complexity of such pipelines and their dependence on external modules limit their practical usability. Researchers have since shifted focus toward investigating the root causes of hallucinations, exploring targeted optimizations in both the decoding strategies and the decoding processes. Some methods Leng et al. (2024); Wang et al. (2024c); Zhang et al. (2025); Zhao et al. (2025) apply contrastive decoding algorithms to refine the logits, thereby reducing the likelihood of hallucination-related tokens. While these approaches improve the expressiveness of the model and enhance the transfer of visual information into textual form, they do not strengthen the model’s intrinsic cross-modal understanding. Other works have proposed enhancements to the attention mechanism, such as correcting visual positional bias Li et al. (2025); Zhu et al. (2025b), directly boosting visual attention Zhu et al. (2025a); He et al. (2025), or redistributing attention weights Kang et al. (2025); Yin et al. (2025). However, these methods primarily concentrate on visual attention and overlook the vision-integrated semantic representation.
@@
In contrast to these approaches, we analyze the information flow between the visual and textual modalities in LVLMs and focus the model’s attention on core information within semantic representations, thereby reducing hallucinations more effectively.
3 Preliminaries
@@
LVLMs consist of a visual encoder and an LLM backbone. High-dimensional features extracted by the visual encoder are concatenated with textual embeddings and jointly fed into the LLM. Subsequently, the visual and semantic features undergo cross-modal interaction through the Self-Attention layers within the LLM. The cross-modal interaction in layer $l$ simplified as $\Delta\mathbf{X}^{l}=\text{softmax}(\mathbf{W})\mathbf{X}^{l}$, where $\mathbf{X}^{l}=[\mathbf{X}^{l}_{V}:\mathbf{X}^{l}_{T}]$ denotes the hidden states of visual features $\mathbf{X}^{l}_{V}$ and semantic features $\mathbf{X}^{l}_{T}$. $\mathbf{W}\in\mathbb{R}^{H\times L\times L}$ is the attention weights, $H$ is the number of attention head and $L$ is the model’s context length.
@@
When generating the (jj+1)-th token, the model’s attention to feature $\mathbf{X}^{l}_{i}$ at layer $l$ across all heads is quantified by the weight value $\mathbf{W}_{i,j}$. A larger $\mathbf{W}_{i,j}$ indicates that the model’s current layer relies more heavily on the information from $\mathbf{X}^{l}_{i}$ when determining the generated content. Therefore, we can intuitively infer that more important features $\mathbf{X}^{l}_{i}$ should generally correspond to larger attention weights $\mathbf{W}_{i,j}$.
@@ 4 Method
Figure 4: The diagram of the two-stage attention distribution alignment. We first categorize attention heads based on their focus on semantic representations into semantic heads and other heads. Semantic heads are further divided into core semantic heads and global semantic heads, depending on whether they attend to core semantic representations. We then align the model’s attention distribution with the core semantic heads through a two-stage optimization process.
@@
We attribute the occurrence of hallucinations to a misalignment between the attention distribution of LVLMs and the visual information flow. To address this, we enhance the visual interpretation capabilities of LVLMs by leveraging semantic representations that encapsulate visual information. We first identify superior attention heads that focus on core semantic representations, and then propagate their strengths to other heads through a two-stage optimization paradigm, as illustrated in Fig. 4.
4.1 Attention Distribution Alignment
@@
Considering that visual information is progressively integrated into semantic representations through forward propagation—and that only a few of these representations aggregate core information—an intuitive approach is to increase the model’s attention to these core semantic representations in the later layers. However, rigidly modifying the model’s attention distribution may disrupt its inherent latent features, potentially resulting in suboptimal performance.
@@
In the multi-head attention mechanism, attention heads naturally diversify after training, forming a heterogeneous and specialized structure Vaswani et al. (2017). Accordingly, we identify superior attention heads that attend to core semantic representations and transfer their attention patterns to other heads, leading to a global refinement of the model’s attention distribution.
Specifically, at each layer, we calculate the attention weight of each head with respect to the semantic representations. We then identify those heads that allocate more than 50% of their attention to semantic representations:
$h\in\left\{\begin{aligned} H_{S}\quad&\text{ if }\sum\mathbf{W}^{h}_{S}>\sum\mathbf{W}^{h}_{O}\\ H_{O}\quad&\text{ if }\sum\mathbf{W}^{h}_{S}\leq\sum\mathbf{W}^{h}_{O}\end{aligned}\right$. (1)
@@
$\mathbf{W}^{h}_{S}$ and $\mathbf{W}^{h}_{S}$ respectively denote the attention weights of head $h$ to semantic and other representations. $H_{S}$ denotes the set of heads that focus more on semantic representations, whereas $H_{O}$ denotes other heads. $H_{S}$ are further categorized based on their attention distributions into Core Semantic Heads ($H_{S_{c}}$), which focus on core semantic representations, and Global Semantic Heads ($H_{S_{g}}$), which fail to effectively attend to them:
@@
$h\in\left\{\begin{aligned} H_{S_{c}}\quad&\text{if}\quad\max\mathbf{W}^{h}_{S}>\kappa\cdot\sum\mathbf{W}^{h}_{S}\\ H_{S_{g}}\quad&\text{if}\quad\max\mathbf{W}^{h}_{S}\leq\kappa\cdot\sum\mathbf{W}^{h}_{S}\end{aligned}\right$. (2)
@@
where $\kappa$ is a hyperparameter and $H_{S_{c}}\cup H_{S_{g}}=H_{S}$ holds. We then take the core semantic heads to construct a target distribution to optimize the model’s attention distribution.
@@
The model’s internal attention is then optimized through a two-stage refinement process.
During this process, we incorporate a weighted smoothing mechanism to adjust the attention distribution of each head:
$\hat{\mathbf{W}}^{h}=\frac{\mathbf{W}^{h}+\omega\mathbf{W}^{\prime}}{1+\omega}$ (3)
where $\hat{\mathbf{W}}^{h}$ denotes the optimized attention weights of head $h$, $\mathbf{W}^{\prime}$ represents the target weight distribution derived from superior heads, and $\omega$ is a hyperparameter that controls the extent of the modification.
We first use the global semantic heads to guide the other heads, aiming to align the model’s attention distribution with the flow of visual information. To this end, we apply average pooling across the global semantic heads to construct the target attention distribution:
$\hat{\mathbf{W}}^{h}_{O}=\frac{\mathbf{W}^{h}_{O}+\omega\cdot\frac{1}{|H_{S_{g}}|}\sum_{t\in H_{S_{g}}}\mathbf{W}^{t}_{S_{g}}}{1+\omega},$ (4)
Subsequently, we use the core semantic heads to guide the global semantic heads, with the goal of enhancing the model’s focus on critical information. Here, we employ max pooling across core semantic heads to form the corresponding target distribution:
$\hat{\mathbf{W}}^{h}_{S_{g}}=\frac{\mathbf{W}^{h}_{S_{g}}+\omega\cdot\max_{t\in H_{S_{c}}}\mathbf{W}^{t}_{S_{c}}}{1+\omega}$ (5)
We provide the algorithm of the two-stage optimization process in Appendix A.
Method LLaVA-1.5 InstructBLIP LLaVA-1.5-13B
@@
Maximum Generation Length is 64
Regular 24.4 8.9 56.6 35.6 13.2 56.4 24.0 7.8 56.5
VCD Leng et al. (2024) 25.0 8.3 59.0 32.2 10.3 60.6 23.4 7.5 59.4
ICD Wang et al. (2024c) 23.2 8.1 58.4 29.8 9.8 60.6 18.8 6.6 58.4
PAI Liu et al. (2024e) 20.0 6.2 56.9 26.0 8.9 53.9 17.4 5.3 58.7
IBD Zhu et al. (2025a) 21.2 6.9 58.8 27.8 9.2 60.3 22.4 7.1 59.7
VAR Kang et al. (2025) 25.2 8.6 55.3 - - - 25.8 8.5 56.3
AD-HH Yang et al. (2025) 19.4 6.3 52.5 - - - 20.0 6.5 53.9
VAF Yin et al. (2025) 26.2 9.3 56.6 32.0 12.0 55.3 23.2 8.0 57.6
DeGF Zhang et al. (2025) 22.4 7.2 58.2 32.4 11.0 59.4 22.4 6.9 59.6
CICD Zhao et al. (2025) 18.0 6.1 59.6 23.8 7.7 62.2 20.4 6.6 59.4
@@
Maximum Generation Length is 512
Regular 54.6 16.4 72.6 62.6 19.5 66.9 58.8 17.0 73.4
VCD Leng et al. (2024) 59.8 17.8 75.6 64.8 18.8 71.9 60.2 16.4 76.9
ICD Wang et al. (2024c) 57.0 15.0 74.6 59.0 17.1 69.2 55.0 14.5 76.4
PAI Liu et al. (2024e) 41.2 10.4 68.6 67.6 19.4 68.0 35.4 9.5 73.1
IBD Zhu et al. (2025a) 57.6 16.5 74.2 57.6 15.7 70.8 50.6 14.3 76.1
VAR Kang et al. (2025) 60.0 18.3 72.6 - - - 54.8 15.3 73.8
AD-HH Yang et al. (2025) 46.6 12.6 66.5 - - - 50.0 13.5 71.0
VAF Yin et al. (2025) 58.8 19.0 71.3 58.6 17.8 66.9 57.2 16.2 73.7
DeGF Zhang et al. (2025) 57.4 16.3 75.5 59.0 17.7 71.5 51.8 14.2 75.5
CICD Zhao et al. (2025) 43.8 11.7 75.0 49.8 13.7 70.3 44.4 12.9 76.0
@@
Table 1: Results on CHAIR. SEVIBalanced and SEVIFocused represent our method operating in the Balanced and Focused mode, respectively. Cs and Ci represent CHAIRs and CHAIRi, R represents Recall.